A and B are complementary angles. If the measure of angle A is increased 9 degrees and the measure of angle B is decreased 15 degrees, then the measure of each of the two angles will be the same. What are the measures of angle A and B?
Answer by checkley77(12844)

A+9=B-15
A+B=90 OR A=90-B
(90-B)+9=B-15
90-B+9=B-15
-B-B=-15-90-9
-2B=-114
B=-114/-2
B=57 DEGREES FOR THE LARGER ANGLE.
A=90-57
A=33 DEGREES FOR THE SMALLER ANGLE.
PROOF:
33+9=57-15
42=42